# Information Density Bounds for Privacy

### v1

@@ II-B Pointwise Maximal Leakage
where $P_{X\mid Y=y}$ denotes the posterior distribution of $X$ given $y\in\mathcal{Y}$. Observe that PML is a context-aware privacy measure since it depends on the prior distribution $P_{X}$.
Viewing $\ell_{P_{XY}}(X\to Y)$ as a function of the random variable $Y$, PML itself becomes a random variable whose various statistical properties can be examined and constrained to provide privacy guarantees. A stringent privacy guarantee can thus be defined by requiring $\ell_{P_{XY}}(X\to Y)$ to be essentially bounded. Formally, we say that a privacy mechanism $P_{Y\mid X}$ satisfies $\varepsilon$-PML if $\ell_{P_{XY}}(X\to Y)\leq\varepsilon$ (almost surely) for some $\varepsilon\geq 0$. Building on this, [62] investigates the design of privacy mechanisms under the $\varepsilon$-PML constraint for secrets with finite alphabets. Specifically, the authors formulate a linear program for maximizing a class of convex utility functions called sub-convex functions [62, Def. 5]. The optimal solutions to this linear program are referred to as PML extremal mechanisms. The paper also introduces the concept of PML privacy regions as a partitioning of the space of the privacy parameter $\varepsilon\in[0,\infty)$ into disjoint intervals. The privacy regions specify an upper bound on the number $(x,y)\in\mathcal{X}\times\mathcal{Y}$ pairs with $P_{Y\mid X=x}(y)=0$. The first region, termed the high-privacy regime, corresponds to the interval $0\leq\varepsilon<\log\frac{1}{1-p_{\mathrm{min}}}$. In this region, no zero probability assignments are allowed in the mechanism. It is then proved that the PML extremal mechanism $P^{*}_{Y\mid X}$ in the high-privacy regime has the following form:
@@
where $i,j\in[\lvert\mathcal{X}\rvert]$ [62, Thm. 3].
@@ II-C Local Information Privacy and Asymmetric Local Information Privacy
Given $\varepsilon_{l},\varepsilon_{u}\geq 0$, a privacy mechanism $P_{Y\mid X}$ is said to satisfy $(\varepsilon_{l},\varepsilon_{u})$-ALIP if $-\varepsilon_{l}\leq i_{P_{XY}}(X;Y)\leq\varepsilon_{u}$ almost surely.
@@ II-D Local Differential Privacy
($\mathbb{P}\left(\left\{\sup_{\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}}\,\log\,\frac{P_{Y\mid X}(\mathcal{E}\mid X)}{P_{Y\mid X}(\mathcal{E}\mid X^{\prime})}\leq\varepsilon\right\}\right)=1,$ (8)
@@
LDP has close connections to context-aware privacy measures. The following result from [18] demonstrates how LDP guarantees translate into PML, LIP, and ALIP guarantees assuming that $\mathcal{X}$ and $\mathcal{Y}$ are finite. Note that while [18, Thm. 1] articulates the LIP result alone, its proof also provides guarantees in terms of ALIP and PML. Recall that $p_{\mathrm{min}}\coloneqq\min\limits_{x\in\mathcal{X}}P_{X}(x)$.
Proposition 1 ([18, Thm. 1]).
@@
If $P_{Y\mid X}$ satisfies $\varepsilon$-LDP, then it also satisfies $\varepsilon_{1}$-LIP, $(\varepsilon_{1},\varepsilon_{2})$-ALIP and $\varepsilon_{2}$-PML.
The previous result is straightforward and intuitive as it illustrates the transition from a context-free property of a mechanism to guarantees that hold for a specific prior distribution. However, the reverse direction is more intricate. It involves moving from a property tied to both a mechanism and a specific prior distribution to a universal property of the mechanism that holds irrespective of the prior. Interestingly, [18, Prop. 1] showed that $\varepsilon$-LIP implies $2\varepsilon$-LDP, and [19, Prop. 1] established that $(\varepsilon_{l},\varepsilon_{u})$-ALIP yields $\varepsilon_{l}+\varepsilon_{u}$-LDP. In general, $\varepsilon$-PML may not provide any LDP guarantee since a mechanism satisfying $\varepsilon$-PML may contain $P_{Y\mid X=x}(y)=0$ for some $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Nonetheless, as we will explore in Section V, in the high-privacy regime PML guarantees indeed translate into valid LDP guarantees.
@@ III Quantifying Information Leaked to Risk-averse Adversaries
The purpose of this section is to explore the implications of defining privacy by imposing a lower bound on the information density $i_{P_{XY}}$. We begin by assuming that $X$ is a finite random variable and demonstrate that given $y\in\mathcal{Y}$, $-\min_{x}i_{P_{XY}}(x;y)$ describes the information leaking to risk-averse adversaries, as opposed to the opportunistic adversaries of PML. In particular, we show that $-\min_{x}i_{P_{XY}}(x;y)$ is operationally meaningful in two adversarial threat models. The first setup called the randomized function model, considers adversaries who wish to minimize the probability of incorrectly guessing the value of an arbitrary (randomized) function of $X$. The second setup, called the cost function model, considers adversaries who wish to minimize arbitrary non-negative cost functions. We refer to the information leakage measure obtained in both threat models as pointwise maximal cost (PMC).
@@ III-A Randomized Function Model
Suppose $\mathcal{X}$ is a finite set. Consider an adversary who attempts to guess the value of a (finite) randomized function of $X$, denoted by $U$. Upon observing $y\in\mathcal{Y}$, the adversary formulates a guess of $U$, denoted by $\hat{U}$ using a kernel $P_{\hat{U}\mid Y}$ that minimizes the probability of an incorrect guess. Essentially, this threat model is an adaptation of [22] and [12, Section II-A] to the risk-averse case, where the adversary seeks to minimize the probability of incorrectly guessing $U$ instead of maximizing the probability of correctly guessing it. To quantify the risk incurred by such an adversary, we compare the adversary’s probability of incorrectly guessing $U$ having access to $y$ and the probability of incorrectly guessing $U$ without access. Consequently, we define
@@
In general, we may not know which function of $X$ the adversary is interested in, or different adversaries may be interested in different functions of $X$. Therefore, to achieve robustness in our model, we define the pointwise maximal cost (PMC) by maximizing $\Lambda_{U}(X\to y)$ over all possible $U$’s satisfying the Markov chain $U-X-Y$.
@@
$-\varepsilon\leq\min_{x\in\mathcal{X}}i(x;y)\iff\Lambda_{P_{XY}}(X\to y)\leq\varepsilon,$ (14)
@@
$\inf_{P_{\hat{U}\mid Y}}\;\mathbb{P}[U\neq\hat{U}\mid Y=y]=1-\sup_{P_{\hat{U}\mid Y}}\;\mathbb{P}[U\neq\hat{U}\mid Y=y]=1-\max_{u\in\mathcal{U}}P_{U\mid Y=y}(u),$
@@
Next, we show that $\Lambda_{P_{XY}}(X\to y)\geq D_{\infty}(P_{X}\|P_{X\mid Y=y})$. Let $x^{*}\in\operatorname*{arg\,max}_{x\in\mathcal{X}}\frac{P_{X}(x)}{P_{X\mid Y=y}(x)}$, and define $V\coloneqq\mathbf{1}_{X\setminus\{x^{*}\}}$. Then, $V$ is a binary random variable with distribution $P_{V}(0)=1-P_{V}(1)=P_{X}(x^{*})$. Let $k$ be a large integer. We define a random variable $W$ with alphabet $\mathcal{W}=\{1,\ldots,k+1\}$ and induced by the kernel
$P_{W\mid V=0}(w)=\begin{cases}\frac{1}{k}&\text{if}\;w\in\{1,\ldots,k\},\\ 0&\text{if}\;w=k+1,\\ \end{cases}$
@@
$P_{W\mid V=1}(w)=\begin{cases}0&\text{if}\;w\in\{1,\ldots,k\},\\ 1&\text{if}\;w=k+1.\\ \end{cases}$
By inspection, $P_{W}(w)=\frac{P_{V}(0)}{k}=\frac{P_{X}(x^{*})}{k}$ for $w\in[k]$ and $P_{W}(k+1)=P_{V}(1)$. Thus, taking $k$ to be sufficiently large, we can ensure that $P_{W}(k+1)=\max_{w}P_{W}(w)$. Moreover, since the Markov chain $W-V-X-Y$ holds $P_{W\mid Y=y}(w)=\frac{P_{V\mid Y=y}(0)}{k}=\frac{P_{X\mid Y=y}(x^{*})}{k}$ for $w\in[k]$ and $P_{W\mid Y=y}(k+1)=P_{V\mid Y=y}(1)=P_{X\mid Y=y}(\mathcal{X}\setminus\{x^{*}\})$. Once again, taking $k$ to be sufficiently large we can ensure that $P_{W\mid Y=y}(k+1)=\max_{w}P_{W\mid Y=y}(y)$.
Since $W-X-Y$ forms a Markov chain, we get
@@
for all $Z$’s satisfying the Markov chain $Z-X-Y$ and all $y\in\mathcal{Y}$. Closedness under pre-processing implies that the amount of information leaking about (random) functions of $X$ can never exceed the amount of information leaking about $X$ itself.
@@ III-B Cost Function Model
Once again, let $\mathcal{X}$ be a finite set. Now, we consider adversaries aiming to minimize the expected value of non-negative cost functions $c:\mathcal{X}\times\mathcal{W}\to\mathbb{R}_{+}$, where $\mathcal{W}$ is an arbitrary finite set representing the adversary’s guessing space. In this scenario, the adversary observes $y\in\mathcal{Y}$ and selects $w\in\mathcal{W}$ to minimize $\mathbb{E}[c(X,w)\mid Y=y]$. Essentially, this threat model is an adaptation of [14] and [12, Section II-B] to the risk-averse case, where the adversary seeks to minimize a cost function instead of maximizing a gain function. To quantify the risk associated with such an adversary, we compare the smallest posterior expected cost with the smallest prior expected cost in a ratio. Accordingly, we define
@@
Similar to the equivalence between the randomized function model and the gain function model established in [12, Thm. 2] for PML, below we demonstrate that (11) and (21) provide equivalent characterizations of PMC. This result is proved in Appendix A.
@@ III-C Relationship to Maximal Cost Leakage and Maximal Realizable Cost
Definition 8 (Maximal cost leakage [22, Def. 11]).
@@
Issa et al. [22, Thm. 15] then showed that maximal cost leakage takes the simple form
@@
Earlier in Remark 2 we argued that PMC is closed under pre-processing; thus, it suffices to consider the costs of guessing $X$ itself. In a similar vein, below we demonstrate that when evaluating the information leaking to the average outcome of $Y$ as in maximal cost leakage, it suffices to analyze the costs associated with guessing $X$ alone. Specifically, we show that the supremum over $U$’s satisfying the Markov chain $U-X-Y$ in (23) is superfluous.
@@
In other words, the supremum over $U$’s satisfying the Markov chain $U-X-Y$ in (23) is superfluous.
@@
Fix a random variable $U$ satisfying the Markov chain $U-X-Y$ and let $\mathcal{U}$ denote the alphabet of $U$. Let $c_{1}:\mathcal{U}\times\mathcal{W}\to\mathbb{R}_{+}$ be a cost function, where $\mathcal{W}$ is a finite set. Consider the cost function $c_{2}:\mathcal{X}\times\mathcal{W}\to\mathbb{R}_{+}$ defined as
@@
where the third equality is due to the Markov chain $U-X-Y$. Hence, for each random variable $U$ satisfying the Markov chain $U-X-Y$ and cost function $c_{1}:\mathcal{U}\times\mathcal{W}\to\mathbb{R}_{+}$, there exists a cost function $c_{2}:\mathcal{X}\times\mathcal{W}\to\mathbb{R}_{+}$ such that
@@
$\mathcal{L}^{\mathsf{c}}(X\to Y)$ $=\sup_{U:U-X-Y}\left(-\log\;\mathbb{E}_{Y\sim P_{Y}}\left[\exp\Big{(}-\Lambda_{P_{UY}}(U\to Y)\Big{)}\right]\right)$ (24a)
$\leq\sup_{U:U-X-Y}\mathbb{E}_{Y\sim P_{Y}}\left[\Lambda_{P_{UY}}(U\to Y)\right]$ (24b)
$\leq\mathbb{E}_{Y\sim P_{Y}}\left[\sup_{U:U-X-Y}\Lambda_{P_{UY}}(U\to Y)\right]$ (24c)
$=\mathbb{E}_{Y\sim P_{Y}}\left[\Lambda_{P_{XY}}(X\to Y)\right],$ (24d)
where (24b) is due to Jensen’s inequality, and (24d) follows from the closedness of PMC under pre-processing. Note that the strict concavity of $\log(\cdot)$ implies that Jensen’s inequality is strict. Consequently, maximal cost leakage does not describe the expected value of PMC but underestimates it.
Definition 9 (Maximal realizable cost [22, Def. 12]).
@@
$\mathcal{L}^{\mathsf{r}\mathsf{c}}(X\to Y)\coloneqq\sup_{\begin{subarray}{c}U:U-X-Y,\\ {\mathcal{W}},c:\mathcal{U}\times\mathcal{W}\to\mathbb{R}_{+}\end{subarray}}\;\log\;\frac{\inf\limits_{w\in\mathcal{W}}\;\mathbb{E}\Big{[}c(U,w)\Big{]}}{\min\limits_{y\in\mathcal{Y}}\;\inf\limits_{\hat{w}\in\mathcal{W}}\;\mathbb{E}\Big{[}c(U,\hat{w})\mid Y=y\Big{]}},$ (25)
@@
Issa et al. [22, Thm. 16] showed that maximal realizable cost can be expressed as

### v2

@@ II-B Pointwise Maximal Leakage
where $P_{X\mid Y=y}$ denotes the posterior distribution of $X$ given $y\in\mathcal{Y}$. Observe that PML is a context-aware privacy measure, that is, it depends on the prior distribution $P_{X}$.
Viewing $\ell_{P_{XY}}(X\to Y)$ as a function of the random variable $Y$, PML itself becomes a random variable whose various statistical properties can be examined and constrained to provide privacy guarantees. A stringent privacy guarantee can thus be defined by requiring $\ell_{P_{XY}}(X\to Y)$ to be essentially bounded. Formally, we say that a privacy mechanism $P_{Y\mid X}$ satisfies $\varepsilon$-PML if $\ell_{P_{XY}}(X\to Y)\leq\varepsilon$ (almost surely) for some $\varepsilon\geq 0$. Building on this, [25] investigates the design of privacy mechanisms under the $\varepsilon$-PML constraint for secrets with finite alphabets. Specifically, the authors formulate a linear program for maximizing a class of convex utility functions called sub-convex functions [25, Def. 5]. The optimal solutions to this linear program are referred to as PML extremal mechanisms. The paper also introduces the concept of PML privacy regions as a partitioning of the space of the privacy parameter $\varepsilon\in[0,\infty)$ into disjoint intervals. The privacy regions specify an upper bound on the number $(x,y)\in\mathcal{X}\times\mathcal{Y}$ pairs with $P_{Y\mid X=x}(y)=0$. The first region, termed the high-privacy regime, corresponds to the interval $0\leq\varepsilon<\log\frac{1}{1-p_{\mathrm{min}}}$. In this region, no zero probability assignments are allowed in the mechanism. It is then proved that the PML extremal mechanism $P^{*}_{Y\mid X}$ in the high-privacy regime has the following form:
@@
where $i,j\in[\lvert\mathcal{X}\rvert$] [25, Thm. 3].
@@ II-C Local Information Privacy and Asymmetric Local Information Privacy
Given $\varepsilon_{l},\varepsilon_{u}\geq 0$, a privacy mechanism $P_{Y\mid X}$ is said to satisfy $(\varepsilon_{l},\varepsilon_{u})$-ALIP if $\varepsilon_{l}\leq i_{P_{XY}}(X;Y)\leq\varepsilon_{u}$ almost surely.
@@ II-D Local Differential Privacy
$\mathbb{P}\left(\left\{\sup_{\mathcal{E}\in\mathcal{S}_{\mathcal{Y}}}\,\log\,\frac{P_{Y\mid X}(\mathcal{E}\mid X)}{P_{Y\mid X}(\mathcal{E}\mid X^{\prime})}\leq\varepsilon\right\}\right)=1,$ (12)
@@
LDP has close connections to context-aware privacy measures. We state two results that illustrate these connections. The first, from [30], demonstrates how LDP guarantees translate into PML, LIP, and ALIP guarantees assuming that $\mathcal{X}$ and $\mathcal{Y}$ are finite given a fixed prior distribution $P_{X}$. Note that while [30, Thm. 1] articulates the LIP result alone, its proof also provides guarantees in terms of ALIP and PML. Recall that $p_{\mathrm{min}}\coloneqq\min\limits_{x\in\mathcal{X}}P_{X}(x)$.
Proposition 1.
@@
If $P_{Y\mid X}$ satisfies $\varepsilon$-LDP, then it also satisfies $\varepsilon_{1}$-LIP, $(\varepsilon_{1},\varepsilon_{2})$-ALIP, and $\varepsilon_{2}$-PML.
The previous result is straightforward and intuitive, as it illustrates the transition from a context-free property of a mechanism to guarantees that hold under a specific prior distribution. However, the reverse direction is more intricate: it involves moving from a property tied to both the mechanism and a prior to a universal property that holds irrespective of the prior. Interestingly, [30, Prop. 1] showed that $\varepsilon$-LIP implies $2\varepsilon$-LDP, and [71, Prop. 1] established that $(\varepsilon_{l},\varepsilon_{u})$-ALIP yields $(\varepsilon_{l}+\varepsilon_{u})$-LDP. In general, $\varepsilon$-PML may not imply any LDP guarantee, since a mechanism satisfying $\varepsilon$-PML may assign $P_{Y\mid X=x}(y)=0$ for some $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Nonetheless, as we will explore in Section V, in the high-privacy regime PML guarantees can indeed translate into valid LDP guarantees.
The second result, from [27] and [22], provides an equivalent formulation of LDP via PML. Although not originally stated in this form, the equivalence follows directly from [27, Thm. 14] and [22, Thm. 3], and leads to the following theorem.
Theorem 1.
Let $\mathcal{P}_{\mathcal{X}}$ denote the set of distributions with full support on $\mathcal{X}$. For any $\varepsilon\geq 0$, the mechanism $P_{Y\mid X}$ satisfies $\varepsilon$-LDP, if and only if it satisfies $\varepsilon$-PML for all $P_{X}\in\mathcal{P}_{\mathcal{X}}$.
We establish a parallel result between LDP and PMC in Section V.
@@ III Quantifying Information Leaked to Risk-averse Adversaries
The purpose of this section is to explore the implications of defining privacy by imposing a lower bound on the information density $i_{P_{XY}}$. We begin by assuming that $X$ is a finite random variable and demonstrate that given $y\in\mathcal{Y}$, $\min_{x}i_{P_{XY}}(x;y)$ describes the information leaking to risk-averse adversaries, as opposed to the opportunistic adversaries of PML. In particular, we show that $\min_{x}i_{P_{XY}}(x;y)$ is operationally meaningful in two adversarial threat models. The first setup called the randomized function model, considers adversaries who wish to minimize the probability of incorrectly guessing the value of arbitrary (randomized) functions of $X$. The second setup, called the cost function model, considers adversaries who wish to minimize arbitrary non-negative cost functions. We refer to the information leakage measure obtained in both threat models as pointwise maximal cost (PMC).
@@ III-A Randomized Function Model
Suppose $\mathcal{X}$ is a finite set. Consider an adversary who attempts to guess the value of a (finite) randomized function of $X$, denoted by $U$. Upon observing $y\in\mathcal{Y}$, the adversary formulates a guess of $U$, denoted by $\hat{U}$ using a kernel $P_{\hat{U}\mid Y}$ that minimizes the probability of an incorrect guess. Essentially, this threat model is an adaptation of [27] and [55, Section II-A] to the risk-averse case, where the adversary seeks to minimize the probability of incorrectly guessing $U$ instead of maximizing the probability of correctly guessing it. To quantify the risk incurred by such an adversary, we compare the probability of incorrectly guessing $U$ with access to $y$ and the probability of incorrectly guessing $U$ without access. Consequently, we define
@@
In general, we may not know which function of $X$ the adversary is interested in, or different adversaries may be interested in different functions of $X$. Therefore, to achieve robustness in our model, we define the pointwise maximal cost (PMC) by maximizing $\Lambda_{U}(X\to y)$ over all possible $U$’s satisfying the Markov chain UXYU-X-Y.
@@
$\varepsilon\leq\min_{x\in\mathcal{X}}i(x;y)\iff\Lambda_{P_{XY}}(X\to y)\leq\varepsilon,$ (18)
@@
and
$\inf_{P_{\hat{U}\mid Y}}\;\mathbb{P}[U\neq\hat{U}\mid Y=y]$ $=1-\sup_{P_{\hat{U}\mid Y}}\;\mathbb{P}[U\neq\hat{U}\mid Y=y]$
$=1-\max_{u\in\mathcal{U}}P_{U\mid Y=y}(u),$
@@
Next, we show that $\Lambda_{P_{XY}}(X\to y)\geq D_{\infty}(P_{X}\|P_{X\mid Y=y})$. Let $x^{*}\in\operatorname*{arg\,max}_{x\in\mathcal{X}}\frac{P_{X}(x)}{P_{X\mid Y=y}(x)}$, and define $V\coloneqq\mathbf{1}_{\mathcal{X}\setminus\{x^{*}\}}$. Then, $V$ is a binary random variable with distribution $P_{V}(0)=1-P_{V}(1)=P_{X}(x^{*})$. Let $k$ be a large integer. We define a random variable $W$ with alphabet $\mathcal{W}=\{1,\ldots,k+1\}$ and induced by the kernel
[…]
@@
[…]
By inspection, $P_{W}(w)=\frac{P_{V}(0)}{k}=\frac{P_{X}(x^{*})}{k}$ for $w\in[k]$ and $P_{W}(k+1)=P_{V}(1)$. Thus, taking $k$ to be sufficiently large, we can ensure that $P_{W}(k+1)=\max_{w}P_{W}(w)$. Moreover, since the Markov chain WVXYW-V-X-Y holds $P_{W\mid Y=y}(w)=\frac{P_{V\mid Y=y}(0)}{k}=\frac{P_{X\mid Y=y}(x^{*})}{k}$ for $w\in[k]$ and $P_{W\mid Y=y}(k+1)=P_{V\mid Y=y}(1)=P_{X\mid Y=y}(\mathcal{X}\setminus\{x^{*}\})$. Once again, taking $k$ to be sufficiently large we can ensure that $P_{W\mid Y=y}(k+1)=\max_{w}P_{W\mid Y=y}(w)$.
Since WXYW-X-Y forms a Markov chain, we get
@@
for all $Z$’s satisfying the Markov chain ZXYZ-X-Y and all $y\in\mathcal{Y}$. Closedness under pre-processing implies that the amount of information leaking about (random) functions of $X$ can never exceed the amount of information leaking about $X$ itself.
@@ III-B Cost Function Model
Once again, let $\mathcal{X}$ be a finite set. Now, we consider adversaries aiming to minimize the expected value of non-negative cost functions $c:\mathcal{X}\times\mathcal{W}\to\mathbb{R}_{+}$, where $\mathcal{W}$ is an arbitrary finite set representing the adversary’s guessing space. In this scenario, the adversary observes $y\in\mathcal{Y}$ and selects $w\in\mathcal{W}$ to minimize $\mathbb{E}[c(X,w)\mid Y=y]$. Essentially, this threat model is an adaptation of [3] and [55, Section II-B] to the risk-averse case, where the adversary seeks to minimize a cost function instead of maximizing a gain function. To quantify the risk associated with such an adversary, we compare the smallest posterior expected cost with the smallest prior expected cost in a ratio. Accordingly, we define
@@
Similar to the equivalence between the randomized function model and the gain function model established in [55, Thm. 2] for PML, below we demonstrate that (15) and (25) provide equivalent characterizations of PMC. This result is proved in Appendix A.
@@ III-C Relationship to Maximal Cost Leakage and Maximal Realizable Cost
Definition 8 (Maximal cost leakage [27, Def. 11]).
@@
Issa et al. [27, Thm. 15] then showed that maximal cost leakage takes the simple form
@@
Earlier in Remark 2 we argued that PMC is closed under pre-processing; thus, it suffices to consider the costs of guessing $X$ itself. In a similar vein, below we demonstrate that when evaluating the information leaking to the average outcome of $Y$ as in maximal cost leakage, it suffices to analyze the costs associated with guessing $X$ alone. Specifically, we show that the supremum over $U$’s satisfying the Markov chain UXYU-X-Y in (27) is superfluous.
@@
In other words, the supremum over $U$’s satisfying the Markov chain UXYU-X-Y in (27) is superfluous.
@@
Fix a random variable $U$ satisfying the Markov chain UXYU-X-Y and let $\mathcal{U}$ denote the alphabet of $U$. Let $c_{1}:\mathcal{U}\times\mathcal{W}\to\mathbb{R}_{+}$ be a cost function, where $\mathcal{W}$ is a finite set. Consider the cost function $c_{2}:\mathcal{X}\times\mathcal{W}\to\mathbb{R}_{+}$ defined as
@@
where the third equality is due to the Markov chain UXYU-X-Y. Hence, for each random variable $U$ satisfying the Markov chain UXYU-X-Y and cost function $c_{1}:\mathcal{U}\times\mathcal{W}\to\mathbb{R}_{+}$, there exists a cost function $c_{2}:\mathcal{X}\times\mathcal{W}\to\mathbb{R}_{+}$ such that
@@
$\mathcal{L}^{\mathsf{c}}(X\to Y)$ (28a)
$=\sup_{U:U-X-Y}\left(-\log\;\mathbb{E}_{Y\sim P_{Y}}\left[\exp\Big(-\Lambda_{P_{UY}}(U\to Y)\Big)\right]\right)$ (28b)
$\leq\sup_{U:U-X-Y}\mathbb{E}_{Y\sim P_{Y}}\left[\Lambda_{P_{UY}}(U\to Y)\right]$ (28c)
$\leq\mathbb{E}_{Y\sim P_{Y}}\left[\sup_{U:U-X-Y}\Lambda_{P_{UY}}(U\to Y)\right]$ (28d)
$=\mathbb{E}_{Y\sim P_{Y}}\left[\Lambda_{P_{XY}}(X\to Y)\right],$ (28e)
where (28c) is due to Jensen’s inequality, and (28e) follows from the closedness of PMC under pre-processing. Note that the strict concavity of $\log(\cdot)$ implies that Jensen’s inequality is strict. Consequently, maximal cost leakage does not describe the expected value of PMC but underestimates it.
Definition 9 (Maximal realizable cost [27, Def. 12]).
@@
$\mathcal{L}^{\mathsf{r}\mathsf{c}}$ $(X\to Y)$ (29)
$\coloneqq\!\!\sup_{\begin{subarray}{c}U:U-X-Y,\\ {\mathcal{W}},c:\mathcal{U}\times\mathcal{W}\to\mathbb{R}_{+}\end{subarray}}\;\log\;\frac{\inf\limits_{w\in\mathcal{W}}\;\mathbb{E}\Big[c(U,w)\Big]}{\min\limits_{y\in\mathcal{Y}}\;\inf\limits_{\hat{w}\in\mathcal{W}}\;\mathbb{E}\Big[c(U,\hat{w})\mid Y=y\Big]},$ (30)
@@
Issa et al. [27, Thm. 16] showed that maximal realizable cost can be expressed as